### v1

# Dynamic Space Filling

Abstract
Dynamic space filling (DSF) is a stochastic process defined on any connected graph. Each vertex can host an arbitrary number of particles forming a pile, with every arriving particle landing on the top of the pile. Particles in a pile, except for the particle at the bottom, can hop to neighboring vertices. Eligible particles hop independently and stochastically, with the overall hopping rate set to unity for every eligible particle at every vertex. When the number of vertices in a graph is equal to the total number of particles, the evolution stops at the moment when every vertex gets occupied by a single particle. We determine the halting time distribution on complete graphs. Using the mapping of the DSF into a two-species annihilation process, we argue that on $d-$dimensional tori with $N\gg 1$ vertices, the average halting time scales with the number of vertices as $N^{4/d}$ when $d\leq 4$ and as $N$ when $d>4$.
I Introduction
Filling space with identical objects is a fascinating subject Aste and Weaire (2008); Torquato and Stillinger (2010). Packings and coverings by balls are especially popular. Balls cannot overlap in packing, but they do overlap in coverings as each point should be covered (i.e., belong to at least one ball). The densest sphere packings and the least dense sphere coverings of $\mathbb{R}^{d}$ are particularly popular subjects Kershner (1939); Rogers (1964); Conway and Sloane (1999); Hales (2005); Schürmann and Vallentin (2005, 2006); Cohn (2016); de Laat and Vallentin (2016); Cohn (2017). The densest sphere packings are known de Laat and Vallentin (2016); Cohn (2017) for $d=1,2,3,8,24$. The least dense sphere coverings are known Kershner (1939); Schürmann and Vallentin (2005, 2006) for $d=1,2$ and conjecturally for $d=24$.
Random sequential adsorption (RSA) is the dynamic counterpart of packing where deposition events leading to the overlap with already present balls are discarded Evans (1993); Talbot et al. (2000); Torquato (2002); Krapivsky et al. (2010). Random sequential covering (RSC) is a dynamic counterpart of covering where deposition events leading to an increase of coverage are accepted Krapivsky (2023); Viot and Krapivsky (2025). In RSA and RSC processes, the evolution stops when the system reaches a jammed state. Jammed states are random, albeit the final jamming density of the infinite space $\mathbb{R}^{d}$ is deterministic. The RSA process stops at the filling fraction $\theta_{\text{RSA}}(d)<1$ depending on the spatial dimension; the RSC process stops at the covering number (the average number of balls covering a point) $\theta_{\text{RSC}}(d)>1$. The filling and covering numbers are analytically known only in one dimension.
We investigate the dynamic space filling (DSF) leading to perfect space filling. The DSF proceeds independently on the connected components of a graph. Therefore, we can limit ourselves to connected graphs. To avoid unnatural behaviors, we consider simple graphs, i.e., graphs without self-loops and multiple edges. We also assume that graphs are regular Diestel (2017), i.e., each vertex has the same number of neighbors. Simple connected $r-$regular graphs with lowest or highest number $r$ of neighbors are easy to classify: The only 2-regular connected graph with $N$ vertices is the ring $R_{N}$, and the only $(N-1)-$regular graph with $N$ vertices is the complete graph $K_{N}$. For $12<r<N-$ 1, the total number of $r-$regular graphs is unknown (rapidly growing with $N$ for any fixed $r$).
The DSF process on a graph with $N$ vertices concerns the dynamics of $N$ random walkers. Random walkers form a pile at each vertex according to the order of their arrival at the vertex: The first arrival is at the bottom, etc. The random walker at the bottom is stuck to the vertex, while other random walkers hop independently and stochastically to neighboring vertices. We set the overall hopping rate to unity, so the hopping rate to each neighbor is $1/r$ in the case of $r-$regular graphs. The DSF process comes to a halt when each vertex is occupied by exactly one random walker.
At first sight, the DSF process looks significantly simpler than the lattice versions of RSA and RSC, such as packing or covering of $\mathbb{Z}^{d}$ with dimers. Indeed, for the RSA and RSC, the final jammed states are understood only in one dimension; for the DSF process, the jammed state is a perfectly filled graph, i.e., it is universal and trivial. However, the approach to the perfect filling is highly non-trivial already in one dimension.
In Sec. II, we analyze the DSF process on complete graphs. We determine the halting time distribution for any $N$. The halting time increases linearly with $N$ and remains a non-self-averaging random variable when $N\to\infty$. In this limit, the halting time distribution approaches the scaling form: $P(T_{N})\to N^{-1}\mathcal{P}(\tau)$ with $\tau=T_{N}/N$. The scaled distribution reads
$\mathcal{P}(\tau)=\sum_{k=-\infty}^{\infty}(-1)^{k+1}\,k^{2}\,e^{-k^{2}\tau}$ (1)
Alternatively, $\mathcal{P}(\tau)$ can be expressed via a theta function.
Apart from the complete graphs, we briefly discuss the DSF on tori $T_{d}(L)=(R_{L})^{d}$, equivalently hypercubes with periodic boundary conditions (Sec. III). Such tori are regular graphs with $N=L^{d}$ and $r=2d$. The DSF process on the infinite lattices $\mathbb{Z}^{d}$ with a density of random walkers equal to unity relaxes to the perfectly filled state. This DSF process admits a mapping into a two-species annihilation process with one species immobile and equal concentrations of both species. The vacuum state of the annihilation process corresponds to the perfect filling of the DSF process. Relying on the asymptotic decay laws for the densities in the two-species annihilation process on the infinite lattices $\mathbb{Z}^{d}$ we estimate a typical halting time on the tori with $N$ vertices:
$T_{N}\sim\begin{cases}N^{4/d}&d<4\\ N&d\geq 4\end{cases}$ (2)
The arguments leading to (2) are heuristic, and hence the estimates (2) are conjectural. Even the nature of the random variable $T_{N}$ on large tori is unsettled when $d<4$; for $d\geq 4$, the halting time $T_{N}$ is a non-self-averaging random variable as we argue in Sec. III.
Apart from the halting time $T_{N}$, the duration $t_{N}$ of the last step, namely, the evolution from a state with $N-2$ perfectly filled vertices to the final perfect filling, is another interesting random quantity. In Sec. IV, we outline how to compute the joint distribution $P(T_{N},t_{N})$ on the complete graphs.
In computing the halting time distribution $P(T_{N})$ and the joint distribution $P(T_{N},t_{N})$ on the complete graphs, we rely on the fact that the evolution of the total number $m$ of empty vertices, $m\to m-1$, is determined by $m$ alone. This crucial property occurs only on complete graphs; generally, the neighbors of each empty vertex and their occupancies matter. One can compute the distributions $P(T_{N})$ and $P(T_{N},t_{N})$, circumventing the detailed knowledge of the evolution of the random variable $m$. However, our methods give the exact Laplace transform of the probability distribution $P_{m}(t)$. In Appendix A, we compute the average and the variance of $m$ on large complete graphs and show that $P_{m}(t)$ becomes Gaussian when $N\gg 1$. In Appendix B, we show how to compute higher cumulants of $m$. The computations quickly become cumbersome, so we compute only the third cumulant.
II DSF on complete graph
The DSF process on a connected graph is isomorphic to the two-species diffusion-controlled annihilation process on the same graph. We denote the particles of the two species by $A$ and $B$ and map the DSF representation into the $A-B$ representation as follows:
  • An empty vertex in the DSF hosts a $B$ particle.
  • A vertex occupied by a single particle in the DSF is empty in the realm of the annihilation process.
  • A vertex occupied by $k+1$ particles in the DSF hosts $k$ particles of type $A$.
Therefore, $A$ and $B$ particles never share the same vertex. The rules of the DSF process imply that non-interacting $A$ particles are undergo independent identical random walks. We call $A$ particles active since they are mobile; passive $B$ particles are immobile.
When an $A$ particle hops into an empty vertex, i.e., the vertex occupied by $B$ particle, the vertex gets perfectly filled, i.e., $A$ and $B$ particles immediately annihilate:
$A+B\to\emptyset$ (3)
In other words, we have a two-species diffusion-controlled annihilation process. We postulate that the number of particles in the DSF process is equal to the number of vertices so that the perfect filling can be achieved. In terms of the annihilation process, this means that the initial numbers of $A$ and $B$ particles are equal; then, the numbers of $A$ and $B$ particles will remain equal forever and the perfect filling corresponds to the vacuum state of the annihilation process.
Consider the DSF process on complete graphs. It is convenient to take the complete graph $K_{N+1}$ so that each active particle hops with rate $1/N$ to every of $N$ neighboring vertices. (Recall that we set the overall hoping rate to unity.) Denote by $[m]$ the state with $m$ empty vertices in the DSF representation. The transition $[m]\to[m-1]$ occurs with rate $r_{m}=m^{2}/N$ as there are $m$ active $A$ particles and $m$ passive $B$ particles. If $t_{m}$ is the transition time from $m$ to $m-1$, the halting time is $T_{N}=t_{1}+t_{2}+\cdots+t_{m_{0}}$, where $m_{0}$ is the initial number of empty vertices. The transition times are exponentially distributed:
$\Pi(t_{m})=r_{m}\,e^{-r_{m}t_{m}}$ (4)
From (4), the average transition time and its variance are given by $\langle t_{m}\rangle=r_{m}^{-1}$ and $\sigma_{m}^{2}=\langle t_{m}^{2}\rangle-\langle t_{m}\rangle^{2}=r_{m}^{-2}$. Therefore the average halting time is
$\langle T_{N}\rangle=\sum_{m=1}^{m_{0}}\langle t_{m}\rangle=N\sum_{m=1}^{m_{0}}m^{-2}$ (5)
The infinite sum is $\sum_{m\geq 1}m^{-2}=\zeta(2)=\frac{\pi^{2}}{6}$. Therefore in the leading order
$\langle T_{N}\rangle=\frac{\pi^{2}}{6}\,N$ (6)
The second moment $\langle T_{N}^{2}\rangle=\langle T_{N}\rangle^{2}+\sum_{1\leq m\leq m_{0}}\sigma^{2}_{m}$ becomes
$\langle T_{N}^{2}\rangle=\langle T_{N}\rangle^{2}+N^{2}\sum_{m=1}^{m_{0}}m^{-4}$ (7)
Recalling that $\sum_{m\geq 1}m^{-4}=\zeta(4)=\frac{\pi^{4}}{90}$ we obtain
$\mu_{2}=\lim_{N\to\infty}\frac{\langle T_{N}^{2}\rangle}{\langle T_{N}\rangle^{2}}=\frac{7}{5}$ (8a)
The initial value $m_{0}\sim N$ and hence replacing the sums in (5) and (7) by infinite sums is exact up to $N^{-1}$ in the first case and up to $N^{-3}$ in the second. The precise value value of $m_{0}$ does not affect the leading behavior. (The random initial distribution gives $m_{0}\approx N/e$, see (23). But the same asymptotic behaviors emerge for the extreme initial condition $m_{0}=N$ describing the evolution starting from all particles initially at a single vertex.)
The normalized moments
$\mu_{p}=\lim_{N\to\infty}\frac{\langle T_{N}^{p}\rangle}{\langle T_{N}\rangle^{p}}$
appear to be rational for all integer positive $p$. Laborious but straightforward calculations give
$\mu_{3}=\frac{93}{35}\,,\quad\mu_{4}=\frac{1143}{175}\,,\quad\mu_{5}=\frac{219}{11}$ (8b)
The rationality of all moments $\mu_{p}$ is proven below.
The moments are non-trivial implying that the halting time is an asymptotically non-self-averaging random variable. Thus for its complete characterization, we must determine the halting time distribution. The halting time distribution $P(T_{N})$ can be extracted from the probability distribution $P_{m}(t)$ of the state of the system. This probability distribution obeys
$\frac{dP_{m}}{dt}=r_{m+1}P_{m+1}-r_{m}P_{m}$ (9)
and the initial condition $P_{m}(0)=\delta_{m,m_{0}}$. The halting time distribution is then found from $P(T_{N})=r_{1}P_{1}(T_{N})$. To treat Eqs. (9) we use the Laplace transform:
[…]
The Laplace transform of the halting time distribution is
[…]
Performing the Laplace transform of Eqs. (9) yields
$(s+r_{m})Q_{m}(s)=r_{m+1}Q_{m+1}(s)$ (12a)
for $11\leq m\leq m_{0}-$ 1 and
$(s+r_{m_{0}})Q_{m_{0}}(s)=1$ (12b)
Starting with (12b) and iterating (12a) we determine $Q_{m}(s)$ for all $m$. In particular, the Laplace transform (11) of the halting time distribution reads
$Q(s)=\prod_{m=1}^{m_{0}}\frac{r_{m}}{r_{m}+s}=\prod_{m=1}^{m_{0}}\left[1+\frac{sN}{m^{2}}\right]^{-1}$ (13)
When $N\to\infty$, the halting time distribution approaches the scaling form
$P(T_{N})\to N^{-1}\,\mathcal{P}(\tau)\quad{\rm with}\quad\tau=\frac{T_{N}}{N}$ (14)
implying that the Laplace transform admits the scaling form
$Q(s)\to\mathcal{Q}(\sigma),\qquad\sigma=sN$ (15)
in the $N\to\infty$ limit. This indeed agrees with Eq. (13) that also gives the scaling form of the Laplace transform
$\mathcal{Q}(\sigma)=\prod_{m\geq 1}\left[1+\frac{\sigma}{m^{2}}\right]^{-1}=\frac{\pi\sqrt{\sigma}}{\sinh\big{(}\pi\sqrt{\sigma}\big{)}}$ (16)
The expansion of $\mathcal{Q}(\sigma)$ near $\sigma=0$ confirms previous results (8a) and (8b) found after laborious calculations. Our empirical observation about the rationality of $\mu_{p}$ also becomes obvious. Expanding $\mathcal{Q}(\sigma)$ to higher orders, one can determine any desirable $\mu_{p}$. For instance,
$\mu_{6}=\frac{12730293}{175175}\,,\quad\mu_{7}=\frac{221157}{715}\,,\quad\mu_{8}=\frac{457141779}{303875}$
Inverting Laplace transform (16) yields the announced distribution (1) of the scaled halting time. The asymptotic behaviors of the scaled halting time distribution are
$\mathcal{P}(\tau)\simeq\begin{cases}\frac{1}{2}\big{(}\frac{\pi}{\tau}\big{)}^{\frac{5}{2}}e^{-\frac{\pi^{2}}{4\tau}}&\tau\to 0\\ 2e^{-\tau}&\tau\to\infty\end{cases}$ (17)
(see also Fig. 1). The large time asymptotic follows from the exact solution (1): When $\tau\to\infty$, the sum on the right-hand side of Eq. (1) is dominated by the terms with $k=\pm 1$. The small time asymptotic is extracted from the $\sigma\to\infty$ behavior of the Laplace transform (16).
Figure 1: The scaled halting time distribution $\mathcal{P}(\tau)$.
During the initial stage far before the halting time, $t\ll N$, the number of $B$ particles is a self-averaging random quantity with average and variance
$\langle m\rangle=Nn,\qquad n=\frac{1}{e+t}$ (18a)
$\langle m^{2}\rangle_{c}=Nv,~{}\quad v=\frac{1}{3}\left[\frac{1}{e+t}+\frac{2e^{3}-3e^{2}}{(e+t)^{4}}\right]$ (18b)
if the initial positions of the particles in the DSF process are uncorrelated. The derivation of Eqs. (18), see Appendix A, takes into account that for the annihilation process $n(0)=e^{-1}$ and $v(0)=e^{-1}(1-e^{-1})$ if the initial positions of the particles in the DSF process are uncorrelated and $N\gg 1$.
The distribution $P_{m}(t)$ acquires the scaling form
$P_{m}(t)=\frac{1}{\sqrt{Nv}}\,\Phi(\xi)$ (19a)
in the scaling region
$N\to\infty,\qquad t\to\infty,\qquad\xi=\frac{m-Nn}{\sqrt{Nv}}=\text{finite}$ (19b)
The scaled distribution is Gaussian with zero mean and unit variance
$\Phi(\xi)=\frac{1}{\sqrt{2\pi}}\,e^{-\xi^{2}/2}$ (19c)
The proof of (18)–(19) is relegated to the Appendix A.
Denote by $F(t_{N})$ the distribution of the duration $t_{N}$ of the final step. This distribution can be found from the same formula $F(t_{N})=r_{1}P_{1}(t_{N})=N^{-1}P_{1}(t_{N})$ as the distribution of the halting time, the difference is that as the initial condition we should use $P_{m}(0)=\delta_{m,1}$ asserting that we start with a single empty vertex. Solving
$\frac{dP_{1}}{dt_{N}}=-\frac{1}{N}\,P_{1},\qquad P_{1}(0)=1$ (20)
yields $F(t_{N})=N^{-1}e^{-\tau}$ with $\tau=t_{N}/N$. Thus, the duration of the last step is a non-self-averaging random variable. The normalized moments are
$\frac{\langle t^{p}\rangle}{\langle t\rangle^{p}}=p!$ (21)
The exponential distribution of the duration of the final time is exact for arbitrary $N$, and hence Eqs. (21) are exact and independent on $N$.
In the case of localized initial condition, namely, when all articles in the DSF process are initially at a single vertex, the initial conditions for the annihilation process are $n(0)=1$ and $v(0)=0$ when $N\gg 1$, and instead of (18) we obtain
$n=\frac{1}{1+t}\,,\qquad v=\frac{1}{3}\left[\frac{1}{1+t}-\frac{1}{(1+t)^{4}}\right]$ (22)
as we show in Appendix A.
III DSF in finite dimensions
We begin with the DSF process on the infinite hyper-cubic lattices $\mathbb{Z}^{d}$. The evolution continues forever. We are mostly interested in the long time behavior. If the initial positions are uncorrelated, the probability $p_{\ell}$ to find $\ell$ particles at a site is
$p_{\ell}=\frac{e^{-1}}{\ell!}$ (23)
We map the DSF process onto a two-species diffusion-controlled annihilation process (3). Our convention that the overall hopping rate of $A$ particles is equal to unity implies that the diffusion coefficient of active particles is $1/(2d)$ on the hypercubic lattice $\mathbb{Z}^{d}$. The initial condition (23) for the DSF process leads to $b(0)=p_{0}$, i.e., the initial densities of $A$ and $B$ particles are
$a(0)=b(0)=e^{-1}$ (24)
The densities of $A$ and $B$ particles remain equal throughout the evolution, $a(t)=b(t)=n(t)$, and decay as
$n(t)\sim\begin{cases}t^{-d/4}&d<4\\ t^{-1}&d\geq 4\end{cases}$ (25)
when $t\gg 1$. A naive ‘mean-field’ treatment suggests that the density obeys $\dot{n}=-n^{2}$ leading to the $t^{-1}$ decay Krapivsky et al. (2010). The decay laws (25) thus assert that the mean-field treatment provides qualitatively correct description above the critical dimension $d\geq d_{c}=4$.
The chief reason for the anomalously slow $t^{-d/4}$ decay is spontaneous ‘phase’ separation that is particularly striking in one dimension. The system spontaneously organizes Ovchinnikov and Zeldovich (1978); Toussaint and Wilczek (1983); Kang and Redner (1984) into a mosaic of alternating $A$ and $B$ domains of typical size $\sqrt{t}$, and the annihilation process (3) occurs at the domain boundaries Leyvraz and Redner (1991); Redner and Leyvraz (1991); Leyvraz and Redner (1992). The width of these reaction regions increases as $t^{3/8}$ in one dimension Leyvraz and Redner (1991); Redner and Leyvraz (1991); Leyvraz and Redner (1992), while the fraction of line covered by reaction regions decreases as $t^{-1/8}$.
The decay laws (25) have been established via heuristic Ovchinnikov and Zeldovich (1978); Toussaint and Wilczek (1983); Kang and Redner (1984) and rigorous Bramson and Lebowitz (1988, 1991) analyses for the symmetric version of the two-species annihilation process (3) when the diffusion coefficients are equal: $D_{A}=D_{B}$. The chief prediction (25) is expected to hold when both diffusion coefficients are positive: $D_{A}\geq D_{B}>0$. Our $B$ particles are immobile. The proof Bramson and Lebowitz (1988, 1991) of the decay laws (25) assumes that both species are mobile, but it seemingly can be extended to the case when one species is immobile.
The spatial organization in the annihilation process with one immobile species was studied mostly in one dimension. The emerging spatial mosaic Leyvraz and Redner (1992) is significantly different from the case when both species have equal diffusivities. Nevertheless, the validity of the decay laws (25) in the extreme case when one species is immobile is supported by simulations Leyvraz and Redner (1992); Sokolov et al. (1991); Reigada et al. (1996) in the case of equal concentarations. (Qualitatively different behaviors emerge in the case of unequal concentrations Sokolov et al. (1991); Reigada et al. (1996).)
Direct numerical simulations of the two-species annihilation is challenging. The mean-field $t^{-1}$ decay holds for a long time in $d=2,3$ before the emergence of the ultimate $t^{-d/4}$ asymptotic. Confirming the $t^{-1/2}$ decay in two dimensions has proven difficult, and the $t^{-3/4}$ decay in three dimensions is essentially impossible to observe. Only the one-dimensional $t^{-1/4}$ decay quickly emerges. A numerical integration of deterministic partial differential equations with random intial densities and reaction term $a({\bf r},t)b({\bf r},t)$ is more amenable in two and three dimensions, and the results support the decay laws (25) for the extreme case of immobile $B$ particles Sokolov et al. (1991); Reigada et al. (1996).
Combining the decay laws (25) and criterion
$Nn(T)\sim 1$ (26)
we arrive at the scaling laws (2) for the halting time $T$. The criterion (26) is difficult to justify as fluctuations might be strong when the expected number of particles $Nn(t)$ is small. To address this issue analytically, it is customary to represent a random variable as a sum of an average that is deterministic and scales linearly with $N$ and a stochastic contribution proportional to $\sqrt{N}$. This so-called Van Kampen expansion Van Kampen (2007) provides an adequate description of many reaction processes, see, e.g., Hendriks et al. (1985); Ben-Avraham and Redner (1986); van Dongen and Ernst (1987); Hilhorst et al. (2004); Boland et al. (2008); Ben-Naim and Krapivsky (2004). In the present situation
$N_{A}=Nn(t)+\sqrt{N}\,\eta$ (27)
The stochastic variable $\eta$ has zero mean, $\langle\eta\rangle=0$, and unknown variance $\langle\eta^{2}\rangle=v(t)$. If the stochastic contribution is sub-dominant, $\sqrt{Nv(t)}\lesssim Nn(t)$ during the evolution process, the naive criterion $Nn(T)\sim 1$ can be used to estimate the halting time. Otherwise, the criterion $\sqrt{Nv(T)}\sim Nn(T)$ is more appropriate. This latter criterion is difficult to justify, and if it works Ben-Naim and Krapivsky (2004); Brilliantov et al. (2021); Otieno et al. (2023); Dyachenko et al. (2023), one should know $v(t)$ to estimate the halting time. The variance $v(t)$ is unknown for the annihilation process (3) in $d$ dimensions. (In Appendix A, we compute $v(t)$ on the complete graphs and show that on complete graphs one can use the naive criterion $Nn(T)\sim 1$.)
Denote by $(1,1)$ the state before halting (a single active $A$ particle and a single passive $B$ particle). Eventually $(1,1)\to(0,0)$ and the evolution stops. Let $\mathcal{T}$ be a typical time between the formation of the $(1,1)$ state and its disapperance. This time, equivalently the time for a random walker to hit a stationary target, scales as
$\mathcal{T}\sim\begin{cases}N^{2}&d=1\\ N\ln(N)&d=2\\ N&d\geq 3\end{cases}$ (28)
Comparing (2) and (28) we see that $T$ greatly exceeds $\mathcal{T}$ in dimensions $d=1,2,3$; in higher dimensions, $d\geq 4$, the times $T$ and $\mathcal{T}$ are comparable. (We digress and note that as usuual, the behavior in two dimensions is most subtle Aldous (1989); Brummelhuis and Hilhorst (1991); Dembo et al. (2004, 2007); Mörters and Peres (2010). Still, a lot is known, e.g., the number of steps it takes for a random walker to completely cover the torus is a self-averaging random variable growing as $CN[\ln N]^{2}$ with known amplitude Dembo et al. (2004).)
The duration $\mathcal{T}$ of the final step before halting is a non-self-averaging random variable. Since $T\sim\mathcal{T}$ when $d\geq 4$, the halting time is a non-self-averaging random variable when $d\geq 4$. The same probably holds also when $d<4$. To gauge this feature quantitatively, one can compare the average $\langle T\rangle$ and the variance $\langle T^{2}\rangle_{c}=\langle T^{2}\rangle-\langle T\rangle^{2}$. The variance is expected to scale algebraically
$\langle T^{2}\rangle_{c}\sim\begin{cases}N^{\beta_{d}}&d<4\\ N^{2}&d\geq 4\end{cases}$ (29)
If $\beta_{d}\geq\frac{4}{d}$ when $d<4$, the halting time is a non-self-averaging random quantity in low spatial dimensions.
IV Discussion
The DSF process on complete graphs (Sec. II) is solvable. We computed the distribution of the halting time and the distribution of the duration of the last step for arbitrary $N$. The former distribution becomes particularly neat in the $N\to\infty$ limit when it approaches to the scaled form (1). One can also probe more convoluted temporal characteristics of the DSF process on complete graphs. For instance, the joint distribution $P(T_{N},t_{N})$ of the duration $T_{N}$ of the process (the halting time) and the duration $t_{N}$ of the last step, $t_{N}<T_{N}$, approaches the scaling form
$P(T_{N},t_{N})=N^{-2}\mathcal{R}(\tau-\tau^{\prime})\,e^{-\tau^{\prime}}$ (30a)
with $(\tau,\tau^{\prime})=N^{-1}(T_{N},t_{N})$. The Laplace transform of the scaled distribution $\mathcal{R}(\tau)$ reads
[…]
When $t_{N}<T_{N}\ll N$, the joint distribution (30) becomes
$P(T_{N},t_{N})\simeq\left(\frac{\pi N}{T_{N}-t_{N}}\right)^{\frac{9}{2}}\,\frac{\exp\!\left[-\frac{\pi^{2}N}{4(T_{N}-t_{N})}-\frac{t_{N}}{N}\right]}{8N^{2}}$
Needless to say, the variables $T_{N}$ and $t_{N}$ are correlated. For instance, from (30a) one finds
[…]
and from (30b) one computes
[…]
implying that
$\frac{\langle T_{N}t_{N}\rangle}{N^{2}}=\frac{\pi^{2}}{6}+1>\frac{\pi^{2}}{6}=\frac{\langle T_{N}\rangle\langle t_{N}\rangle}{N^{2}}$ (32)
A similar calculation yields
$\frac{\langle T_{N}t_{N}^{p}\rangle}{N^{p+1}}=p!\,\frac{\pi^{2}}{6}+p\cdot p!>p!\,\frac{\pi^{2}}{6}=\frac{\langle T_{N}\rangle\langle t_{N}^{p}\rangle}{N^{p+1}}$ (33)
To understand the DSF process in finite dimensions (Sec. III) we mapped it into two-species annihilation process with exactly equal numbers of particles of both species and particles of one species performing nearest-neighbor random walk and immobile particles of another species. This mapping leads to the decay laws (25) for the densities of the species which we combined with naive criterion (26) to estimate the halting time (2). More work is required to justify or disprove (2) in $d<4$.
Conjecturally, the halting time remains a non-self-averaging random variable even in the $N\to\infty$ limit. One can justify this assertion when $d\geq 4$ by noting that (i) the duration of the final step is a non-self-averaging random variable, and (ii) the duration of the final step is comparable with halting time when $d\geq 4$. If $d\leq 3$, the duration of the final step is asymptotically negligible compared to the halting time, so we cannot use the same argument as in $d\geq 4$.
An extension to regular random graphs Wormald (1999) is an interesting avenue for future work. A regular random graph with degree $r\geq 3$ picked uniformly from all possible $r-$regular graphs is effectively infinite-dimensional as manifested by a logarithmic growth of the diameter with $N$. The halting time is expected to be non-self-averaging. The distribution of the halting time should approach a scaling form (14), and only the scaled distribution probably depends on $r$. One can also consider planar regular random graphs (see Di Francesco et al. (1995); Bouttier et al. (2002) and references therein). The remarkable feature of planar regular random graphs is that their intrinsic dimension is $D=4$. Several equilibrium characteristics of statistical physics models on planar regular random graphs are more tractable than on the regular $2D$ lattices. Non-equilibrium processes on planar regular random graphs remain unexplored.
Appendix A Derivation of (18)–(19) and (22)
We assume the linear in $N$ scaling of the average and variance, $\langle m\rangle=Nn(t)$ and $\langle m^{2}\rangle_{c}=Nv(t)$, of the number of empty vertices. The probability distribution $P_{m}(t)$ satisfies (9) with $r_{m}=m^{2}/N$, i.e.,
$N\,\frac{dP_{m}}{dt}=(m+1)^{2}P_{m+1}-m^{2}P_{m}$ (34)
We are interested in the regime where $m\sim N$, so we treat $m$ as a continuous variable and reduce a set of ordinary differential equations (9) into a single partial differential equation
$\partial_{t}P=N^{-1}\left(\partial_{m}+\frac{1}{2}\partial_{m}^{2}\right)\left(m^{2}P\right)$ (35)
Using (19b) we compute the derivatives (dot denotes the derivative with respect to $t$)
$\partial_{m}=\frac{\partial_{\xi}}{\sqrt{Nv}}\,,~{}\partial_{m}^{2}=\frac{\partial_{\xi}^{2}}{Nv}\,,~{}\partial_{t}=-\left[\sqrt{N}\,\frac{\dot{n}}{\sqrt{v}}+\frac{\xi\dot{v}}{2v}\right]\partial_{\xi}$
which we substitute into (35) and obtain
$\begin{split}&-\sqrt{N}\,\frac{\dot{n}}{\sqrt{v}}\,\partial_{\xi}\Phi-\frac{\dot{v}}{2v}\,\partial_{\xi}(\xi\Phi)\\ &=\left(\frac{\partial_{\xi}}{\sqrt{Nv}}+\frac{\partial_{\xi}^{2}}{2Nv}\right)\left[\big{(}\sqrt{N}\,n+\xi\sqrt{v}\big{)}^{2}\Phi\right]\end{split}$ (36)
where we have used the scaling form (19a). Equating the leading $O(\sqrt{N})$ terms in (36) we arrive at a ‘mean-field’ equation for the density of empty vertices:
$\dot{n}=-n^{2}$ (37)
Equating the subleading $O(1)$ terms in (36) yields
$n^{2}\partial_{\xi}^{2}\Phi+(\dot{v}+4nv)\partial_{\xi}(\xi\Phi)=0$ (38)
Integrating (38) gives $\Phi=\exp\!\left[-\frac{\dot{v}+4nv}{2n^{2}}\xi^{2}\right]$ up to the normalization factor. Since the variable $\xi$ has zero mean and unit variance, $\Phi(\xi)$ must be the normal distribution (19c), and therefore
$\dot{v}+4nv=n^{2}$ (39)
Dividing (39) by (37) gives $\frac{dv}{dn}=\frac{4v}{n}-1$, from which $v=Cn^{4}+n/3$. Solving (37) and using initial conditions to fix constant $C$ we finally arrive at
$n=\frac{n_{0}}{1+n_{0}t}\,,\quad v=\frac{n}{3}+\left(v_{0}-\frac{n_{0}}{3}\right)\left(\frac{n}{n_{0}}\right)^{4}$ (40)
The probability that the evolution begins with $m$ empty vertices is $\binom{N}{m}e^{-m}\big{(}1-e^{-1}\big{)}^{N-m}$ for the uncorrelated initial condition. For this binomial distribution we find $n_{0}=e^{-1}$ and $v_{0}=e^{-1}(1-e^{-1})$, so Eqs. (40) reduce to the announced solution (18).
If all particles are initially at a single vertex of the complete graph, we have $n_{0}=1$ and $v_{0}=0$, so Eqs. (40) reduce to the announced solution (22). The distribution $P_{m}(t)$ approaches a Gaussian form, Eq. (19c), in the scaling region (19b) with $n$ and $v$ given by (22). The exact solution of (9) subject to the initial condition $P_{m}(0)=\delta_{m,N}$ is also simple: The Laplace transform of $P_{m}(t)$ defined by (10) reads
$Q_{m}(s)=\frac{N}{m^{2}}\prod_{\ell=m}^{N}\left[1+\frac{sN}{\ell^{2}}\right]^{-1}$ (41)
The ratio of the variance of a random quanity to its average is known as a Fano factor Fano (1947). Equations (40) show that the Fano factor associated with the number of empty verticies quickly approaches a universal (independent on the initial condition) value:
$\frac{\langle m^{2}\rangle_{c}}{\langle m\rangle}=\frac{1}{3}$ (42)
for $t\gg 1$. More precisely, (42) is valid sufficiently far from the halting time, i.e., when $1\ll t\ll N$.
Appendix B Higher cumulants
Here we first present an alternative derivation of (37) and (39). We then show how to probe higher cumulants, and fully determine the third cumulant. The governing equations for the moments $\langle m^{k}\rangle=\sum m^{k}P_{m}(t)$
$N\,\frac{d\langle m\rangle}{dt}$ $=-\langle m^{2}\rangle$ (43a)
$N\,\frac{d\langle m^{2}\rangle}{dt}$ $=-2\langle m^{3}\rangle+\langle m^{2}\rangle$ (43b)
$N\,\frac{d\langle m^{3}\rangle}{dt}$ $=-3\langle m^{4}\rangle+3\langle m^{3}\rangle-\langle m^{2}\rangle$ (43c)
etc. immediately follow from Eqs. (34).
Equation (43a) for the first moment leads to (37) in the leading order. Subtracting from (43b) equation (43a) multiplied by $2\langle m\rangle$ we obtain the evolution equation for the variance $\langle m^{2}\rangle_{c}=\langle m^{2}\rangle-\langle m\rangle^{2}$:
$N\,\frac{d\langle m^{2}\rangle_{c}}{dt}=-2\langle m^{3}\rangle+\langle m^{2}\rangle+2\langle m^{2}\rangle\langle m\rangle$ (44)
Recalling that
$\langle m^{3}\rangle=\langle m\rangle^{3}+3\langle m\rangle\langle m^{2}\rangle_{c}+\langle m^{3}\rangle_{c}$ (45)
we re-write (44) as
$N\,\frac{d\langle m^{2}\rangle_{c}}{dt}$ $=$ $\langle m\rangle^{2}-4\langle m\rangle\langle m^{2}\rangle_{c}$ (46)
$+$ $\langle m^{2}\rangle_{c}-2\langle m^{3}\rangle_{c}$
Similarly to the average and the variance, the third cumulant scales linearly, $\langle m^{3}\rangle_{c}=Nw(t)$. Therefore only the terms in the top layer of (46) are asymptotically relevant, and we verify that (46) reduces to (39) in the leading order.
Similarly we recast Eq. (43c) into an equation for the third cumulant (45):
$N\,\frac{d\langle m^{3}\rangle_{c}}{dt}$ $=$ $-6\langle m^{3}\rangle_{c}\langle m\rangle-\langle m\rangle^{2}$ (47)
$+$ $6\langle m^{2}\rangle_{c}\left[\langle m\rangle-\langle m^{2}\rangle_{c}\right]$
We kept only the leading terms in the right-hand side of (47). In calculations leading to (47) we used (45) and
$\langle m^{4}\rangle$ $=$ $\langle m\rangle^{4}+6\langle m\rangle^{2}\langle m^{2}\rangle_{c}+3\langle m^{2}\rangle_{c}^{2}$ (48)
$+$ $4\langle m^{3}\rangle_{c}\langle m\rangle+\langle m^{4}\rangle_{c}$
Equation (47) yields
$\dot{w}+6nw=-n^{2}+6v(n-v)$ (49)
Dividing (49) by (37) gives
$\frac{dw}{dn}=\frac{6w}{n}+1-\frac{6v}{n}\left(1-\frac{v}{n}\right)$ (50)
which has a simple polynomial solution
$w=\frac{n}{15}+an^{4}+bn^{6}+cn^{7}$ (51)
with amplitudes $a,b,c$ determined by the initial values $n_{0},v_{0},w_{0}$. In the extreme case when all particles are initially at a single vertex of the complete graph, we have $n_{0}=1$ and $v_{0}=w_{0}=0$, and (51) becomes
$w=\frac{n-5n^{4}-6n^{6}+10n^{7}}{15}$ (52)
The ratio of the third cumulant to the averge is another Fano factor which quickly approaches a universal (independent on the initial condition) value:
$\frac{\langle m^{3}\rangle_{c}}{\langle m\rangle}=\frac{1}{15}$ (53)
for $t\gg 1$.
The Fano factors (42) and (53) appear in numerous quantum and classical problems Lee et al. (1995); Blanter and Büttiker (2000); Bodineau and Derrida (2004). If the Fano factors associated with the random variable $m$ are the same as the Fano factors in those problems, the next Fano factor should be $\frac{\langle m^{4}\rangle_{c}}{\langle m\rangle}=-\frac{1}{105}$. It could be possible to verify or disprove this conjectural value by extending the calculations presented in this Appendix. Such pedestrian calculations quickly become unwieldy, so a more sophisticated approach is necessary to deduce the cumulant generating function.